# Fast Ewald Summation using Prolate Spheroidal Wave Functions

### v1

@@ Abstract
Fast Ewald summation efficiently evaluates Coulomb interactions and is widely used in molecular dynamics simulations. It is based on a split into a short-range and a long-range part, where evaluation of the latter is accelerated using the fast Fourier transform (FFT). The accuracy and computational cost depend critically on the mollifier in the Ewald split and the window function used in the spreading and interpolation steps that enable the use of the FFT. The first prolate spheroidal wavefunction (PSWF) has optimal concentration in real and Fourier space simultaneously, and is used when defining both a mollifier and a window function. We provide a complete description of the method and derive rigorous error estimates. In addition, we obtain closed-form approximations of the Fourier truncation and aliasing errors, yielding explicit parameter choices for the achieved error to closely match the prescribed tolerance. Numerical experiments confirm the analysis: PSWF-based Ewald summation achieves a given accuracy with significantly fewer Fourier modes and smaller window supports than Gaussian- and B-spline-based approaches, providing a superior alternative to existing Ewald methods for particle simulations.
@@ 1 Introduction
Long-range interactions such as Coulomb and Stokes potentials arising in particle-based systems are central to molecular dynamics, fluid dynamics, and wave propagation. A major computational challenge is that direct evaluation of such interactions scales quadratically with the number of particles. Ewald summation, introduced in 1921 [12] reduces this cost and has since become a standard tool in large-scale simulations.
In Ewald’s original formulation, the three-dimensional Laplace kernel $G(r)=1/r$ ($r=|\bm{x}|$, $\bm{x}\in\mathbb{R}^{3}$) is decomposed into a mollified kernel $M(r)$ and a residual kernel $R(r)$ according to
@@
where $\operatorname{erf}(x)=\tfrac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}\,\mathrm{d}t$ and $\operatorname{erfc}(x)=1-\operatorname{erf}(x)$ are the error and complementary error functions. The parameter $\sigma>0$ is a width-parameter of the mollification that controls the decay of both terms. In the original Ewald-split (1), the mollified kernel $M$ is the convolution of $G$ with a Gaussian $\tfrac{1}{\sqrt{\pi}\sigma}e^{-r^{2}/\sigma^{2}}$, while the residual is R=GMR=G-M. Because a Gaussian is not compactly supported, $R$ must be truncated numerically outside a cutoff radius $r_{c}>0$, with truncation error controlled by a numerical tolerance that is typically denoted by $\varepsilon$.
@@
where $\{\rho_{j}\}_{j=1}^{n}$ are particle strengths (e.g., charges), $\bm{r}\in\mathbb{Z}^{3}$ indexes the periodic images, and the prime indicates that the self-interaction ($i=j$, $\bm{r}=0$) is omitted. The Ewald split (1) makes the otherwise conditionally convergent sum (2) convergent provided that the charge-neutrality condition $\sum_{j=1}^{n}\rho_{j}=0$ holds. In (3), the mollified part $M$ is evaluated in Fourier space, where $\hat{M}(\bm{k})$ decays rapidly, while the residual $R$ captures local interactions within a cutoff radius $r_{c}$. This reduces the cost of evaluating $\phi$ from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n^{3/2})$ for close to randomly uniform particle distributions [15].
@@
Several fast Ewald variants accelerate the Fourier-space sum further to $\mathcal{O}(n\log n)$ using the fast Fourier transform (FFT) [7]. These include the Particle Mesh Ewald (PME) [8], Smooth Particle Mesh Ewald (SPME) [11], Particle–Particle–Particle–Mesh Ewald (PPPM or P3M) [15], the Spectral Ewald (SE) [20, 28, 2], and the Particle–Particle NFFT Ewald (P2NFFT) method [22, 16]. These schemes spread particle data onto a uniform grid using a localized window function, apply the FFT, scale in Fourier space, and interpolate back. The accuracy and efficiency therefore depend on two distinct choices of functions: (i) the mollifier used in the Ewald split, and (ii) the window function used in spreading and interpolation. Both influence the spectral decay and thus the number of Fourier modes and grid points required for a target accuracy. Classical methods differ mainly in these choices: PME, SPME, and P3M employ low-order B-splines as window functions, while SE and P2NFFT use Gaussians or various types of approximations to the first prolate spheroidal wavefunction.
@@
The first prolate spheroidal wavefunction of order zero (PSWF), studied extensively by Slepian and collaborators [31, 30] and in [24, 23], is optimally concentrated in both real and Fourier space. This makes it attractive both as a mollifier (for splitting) and as a window function (for spreading/interpolation), though in principle different PSWFs with different parameters may be used for the two roles. Historically, the absence of a closed form limited practical use, and approximations such as the Kaiser–Bessel and exponential of semicircle functions were employed instead [5, 18], including in SE and P2NFFT implementations [28, 22]. With modern evaluation algorithms [23] and polynomial approximations [5], PSWFs have become computationally practical.
@@
Until recently, PSWFs have primarily been considered as window functions in fast Ewald summation. However, employing them also as mollifiers in the Ewald split yields further advantages, reducing the number of required Fourier modes by around a factor of two in each dimension. Figure 1 illustrates this effect by comparing the decay of the PSWF and Gaussian functions.
@@
The idea of using PSWFs for kernel splitting was first introduced in the Dual-split Multilevel Kernel (DMK) framework [17]. The PSWF split and the DMK algorithm have since also been applied to Stokes potentials [1]. While the DMK framework established the basic concept, a complete mathematical description of PSWF-based Ewald methods is still lacking.
In this paper, we provide such a description, incorporating the first PSWF of order zero both as a mollifier and a window function, with distinct parameters for the two roles. Our main contributions are:
@@
  • Sharp error bounds for both splitting and windowing, with simplified closed-form approximations for practical use.
  • Guidelines for parameter selection for the Fourier sum, enabling predictive accuracy control without ad-hoc tuning.
@@
Recent work by Liang et al. [19] has emphasized high-performance implementation and benchmarking. Taken together, the results presented here and those of Liang et al. highlight both the theoretical foundations and the practical advantages of PSWF-based Ewald summation.
@@ 2.1 Basic definitions and notation
Fourier-space variables are $\bm{k}\in\mathbb{Z}^{3}$ and $\bm{\omega}:=(2\pi/L)\bm{k}\in\mathbb{R}^{3}$ with $\omega=|\bm{\omega}|$. For a function $f(\bm{x})$, its Fourier transform is $\hat{f}(\bm{\omega})$ (or $\hat{f}(\bm{k})$ when the scaling is clear), Fourier-series coefficients are $c_{\bm{k}}(f)$, and discrete Fourier transform coefficients are denoted $\hat{f}_{\bm{k}}\approx c_{\bm{k}}(f)$. Radially symmetric functions are written $f(r)$ with $r=|\bm{x}|$.
@@ 2.2 A fundamental lemma for Laplacian kernel splitting
with $g(\bm{0})$ defined by continuity. Then
$\hat{g}(\bm{\omega})\;=\;\frac{4\pi}{|\bm{\omega}|^{2}}\,\hat{f}(|\bm{\omega}|).$ (6)
@@
Let $\eta\colon\mathbb{R}^{3}\to\mathbb{R}$ be the radial extension of $f$ (see Appendix C.1), i.e. $\hat{\eta}(\bm{\omega})=\hat{f}(|\bm{\omega}|)$. Thus $\hat{g}(\bm{\omega})=\hat{G}(\bm{\omega})\,\hat{\eta}(\bm{\omega})$ with $\hat{G}(\bm{\omega})=4\pi/|\bm{\omega}|^{2}$, i.e. $g=G*\eta$.
@@ 2.4 Window functions, periodization and Fourier coefficients
On the periodic domain $\Omega$, a periodized window function $\tilde{\varphi}$ is employed. More generally, for any $f$ on $\mathbb{R}^{3}$ we define the periodization $\tilde{f}$ on $\Omega$ by
@@
with coefficients (see Appendix C.2 for the derivation)
@@ 2.5 Prolate Spheroidal Wave Functions
The prolate spheroidal wave functions (PSWFs), $\{\psi_{j}^{c}\}_{j=0}^{\infty}$, arise as eigenfunctions of the integral operator
@@
$\lambda_{j}\psi_{j}^{c}(s)=\int_{-1}^{1}\psi_{j}^{c}(t)\,e^{\mathrm{i}cst}\,dt,$ (23)
with eigenvalues ordered by decreasing modulus
@@
The functions $\psi_{j}^{c}$ are real, orthonormal in $L^{2}([-1,1])$, and alternate parity with $j$ [24].
@@
Thus only $\psi_{0}^{c}$ is needed in the present work. Its Fourier transform satisfies a simple relation:
@@
Let $\hat{\psi}_{0}^{c}(\omega)=\int_{-1}^{1}\psi_{0}^{c}(x)e^{\mathrm{i}\omega x}\,dx$. For $\omega\in[-c,c]$,
$\hat{\psi}_{0}^{c}(\omega)=\lambda_{0}\,\psi_{0}^{c}(\omega/c).$ (27)
For computation, $\psi_{0}^{c}$ can be obtained from its Legendre expansion and a tridiagonal eigenproblem derived from the differential equation [23]
$\frac{\mathrm{d}}{\mathrm{d}x}\!\Big((1-x^{2})\frac{\mathrm{d}\psi_{0}^{c}}{\mathrm{d}x}\Big)+(\chi_{0}-c^{2}x^{2})\psi_{0}^{c}=0,$ (28)
with details given in Appendix B.
@@ 3 Ewald summation
Before presenting the PSWF-based implementation, we briefly review the classical Ewald summation and its components. We then summarize the general fast Ewald algorithm, in which the PSWF will serve as both mollifier and window.
@@ 3.1 Ewald summation for triply periodic problems
Recall, the Ewald split of a radial kernel $G\colon\mathbb{R}^{3}\to\mathbb{R}$ is
@@ 3.1.2 Direct computation of the Fourier-space sum
$\phi_{i}^{\text{far}}\approx\frac{1}{V}\sum_{\bm{k}\in\mathcal{I}_{\bm{m}}\setminus\{\bm{0}\}}\hat{M}(\bm{k})\,\hat{\rho}(\bm{k})\,e^{-\frac{2\pi}{L}\mathrm{i}\,\bm{k}\cdot\bm{x}_{i}},\qquad i=1,\ldots,n.$ (37)
Here,
$\hat{M}(\bm{k})$ $=\hat{G}\bigl(|\bm{\omega}(\bm{k})|\bigr)\,\hat{\gamma}_{\sigma}\bigl(|\bm{\omega}(\bm{k})|\bigr)=\frac{L^{2}}{\pi|\bm{k}|^{2}}\,\hat{\gamma}_{\sigma}\bigl(|\bm{\omega}(\bm{k})|\bigr),\qquad\bm{\omega}(\bm{k}):=\tfrac{2\pi}{L}\bm{k},$ (38)
@@
$\hat{\rho}(\bm{k}):=\sum_{j=1}^{n}\rho_{j}\,e^{\tfrac{2\pi}{L}\mathrm{i}\,\bm{k}\cdot\bm{x}_{j}},\qquad\bm{k}\in\mathcal{I}_{\bm{m}}.$

### v2

@@ Abstract
Fast Ewald summation efficiently evaluates Coulomb interactions and is widely used in molecular dynamics simulations. It is based on a split into a short-range and a long-range part, where evaluation of the latter is accelerated using the fast Fourier transform (FFT). The accuracy and computational cost depend critically on the mollifier in the kernel split and the window function used in the spreading and interpolation steps that enable the use of the FFT. The first prolate spheroidal wavefunction (PSWF) has optimal concentration in real and Fourier space simultaneously, and is used when defining both a mollifier and a window function. We provide a complete description of the method and derive rigorous error estimates. In addition, we obtain closed-form approximations of the Fourier truncation and aliasing errors, yielding explicit parameter choices for the achieved error to closely match the prescribed tolerance. Numerical experiments confirm the analysis: PSWF-based Ewald summation achieves a given accuracy with significantly fewer Fourier modes and smaller window supports than Gaussian- and B-spline-based approaches, providing a superior alternative to existing Ewald methods for particle simulations.
@@ 1 Introduction
Long-range interactions such as Coulomb and Stokes potentials arising in particle-based systems are central to molecular dynamics (MD), fluid dynamics, and wave propagation. A major computational challenge is that direct evaluation of such interactions scales quadratically with the number of particles. Ewald summation, introduced in 1921 [12] reduces this cost and has since become a standard tool in large-scale simulations.
In Ewald’s original formulation, the three-dimensional Laplace kernel $G(r)=1/r$ ($r=|\bm{x}|$, $\bm{x}\in\mathbb{R}^{3}$) is split into a mollified kernel $M(r)$ and a residual kernel $R(r)$ according to
@@
where $\operatorname{erf}(x)=\tfrac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}\,\mathrm{d}t$ and $\operatorname{erfc}(x)=1-\operatorname{erf}(x)$ are the error and complementary error functions. The parameter $\sigma>0$ is a width-parameter of the mollification that controls the decay of both terms. In Ewald’s original kernel split (1), the mollified kernel $M$ is the convolution of $G$ with a Gaussian $\tfrac{1}{\sqrt{\pi}\sigma}e^{-r^{2}/\sigma^{2}}$, while the residual is R=GMR=G-M. Because a Gaussian is not compactly supported, $R$ must be truncated numerically outside a cutoff radius $r_{c}>0$, with truncation error controlled by a numerical tolerance that is typically denoted by $\varepsilon$.
@@
where $\{\rho_{j}\}_{j=1}^{n}$ are particle strengths (e.g., charges), $\bm{r}\in\mathbb{Z}^{3}$ indexes the periodic images, and the prime indicates that the self-interaction ($i=j$, $\bm{r}=0$) is omitted. The Ewald split (1) makes the otherwise conditionally convergent sum (2) convergent provided that the charge-neutrality condition $\sum_{j=1}^{n}\rho_{j}=0$ holds. In (3), the mollified part $M$ is evaluated in Fourier space, where $\hat{M}(\bm{k})$ decays rapidly, while the residual $R$ captures local interactions within a cutoff radius $r_{c}$. This reduces the computational cost of evaluating $\phi$ from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n^{3/2})$ for close to randomly uniform particle distributions [15].
@@
Several fast Ewald variants accelerate the Fourier-space sum further to $\mathcal{O}(n\log n)$ using the fast Fourier transform (FFT) [7]. These include the Particle Mesh Ewald (PME) [8], Smooth Particle Mesh Ewald (SPME) [11], Particle–Particle–Particle–Mesh Ewald (PPPM or P3M) [15], the Spectral Ewald (SE) [20, 28, 2], and the Particle–Particle NFFT Ewald (P2NFFT) method [21, 16]. These schemes spread particle data onto a uniform grid using a localized window function, apply the FFT, scale in Fourier space, and interpolate back. The accuracy and efficiency therefore depend on two distinct choices of functions: (i) the mollifier used in the kernel split, and (ii) the window function used in spreading and interpolation. Both influence the spectral decay and thus the number of Fourier modes and grid points required for a target accuracy. Classical methods differ mainly in these choices: PME, SPME, and P3M employ low-order B-splines as window functions, while SE and P2NFFT use Gaussians or various types of approximations to the first prolate spheroidal wavefunction of order zero (PSWF).
@@
The PSWF, studied extensively by Slepian and collaborators [29, 30] and in [23, 24], is optimally concentrated in both real and Fourier space. This makes it attractive both as a mollifier in the kernel split and as a window function for spreading/interpolation, though in principle different PSWFs with different parameters may be used for the two roles. Historically, the absence of a closed form limited practical use, and approximations such as the Kaiser–Bessel (KB) and exponential of semicircle (ES) functions were employed instead [5, 18], including in SE and P2NFFT implementations [28, 21]. With modern evaluation algorithms [24] and polynomial approximations [5], PSWFs have become computationally practical.
@@
Until recently, PSWF analogues, such as the KB and ES functions, have primarily been considered as window functions in fast Ewald summation. However, employing them also as mollifiers in the kernel split yields further advantages, reducing the number of required Fourier modes by around a factor of two in each dimension compared to Gaussians, as illustrated in Figure 1. PSWF-based kernel splitting was first introduced for the Laplace and Yukawa kernels in the Dual-space Multilevel Kernel-splitting framework (DMK) [17], and was later extended to kernels arising in Stokes flow [1].
@@
Fast Ewald summation based on this type of kernel splitting was developed independently in the present work and by Liang et al. [19]. In their work, a PSWF-based fast Ewald summation method for Coulomb interactions in molecular dynamics is introduced, in which PSWFs are used both for the kernel split and for spreading/interpolation, and the resulting method is integrated into the MD codes GROMACS [35] and LAMMPS [34]. Their benchmarks show substantial practical speedups, with total computation times reduced by up to a factor of three when running large-scale simulations with thousands of cores.
The present work is complementary to [19]. Our focus is on the mathematical formulation and analysis of PSWF-based Ewald summation, including:
@@
  • Rigorous error estimates for both splitting and windowing, with simplified closed-form approximations for practical use.
  • Guidelines for parameter selection for the Fourier sum, enabling predictive accuracy control without ad hoc tuning.
@@ 2.1 Basic definitions and notation
Fourier-space variables are $\bm{k}\in\mathbb{Z}^{3}$ and $\bm{\omega}:=(2\pi/L)\bm{k}\in\mathbb{R}^{3}$ with $\omega=|\bm{\omega}|$. We use the Fourier transform convention $\hat{f}(\bm{\omega})=\int_{\mathbb{R}^{3}}f(\bm{x})\,e^{-\mathrm{i}\bm{\omega}\cdot\bm{x}}\,d\bm{x}$. When convenient, we write $\hat{f}(\bm{k}):=\hat{f}\!\left((2\pi/L)\bm{k}\right)$. Fourier-series coefficients are $c_{\bm{k}}(f)$, and discrete Fourier transform coefficients are denoted $\hat{f}_{\bm{k}}\approx c_{\bm{k}}(f)$. Radially symmetric functions are written $f(r)$ with $r=|\bm{x}|$.
@@ 2.2 A fundamental lemma for Laplacian kernel splitting
with $g(\bm{0})$ defined by continuity. Then, with Fourier transform taken in $\mathbb{R}^{3}$,
$\hat{g}(\bm{\omega})\;=\;\frac{4\pi}{|\bm{\omega}|^{2}}\,\hat{f}(|\bm{\omega}|),\qquad\bm{\omega}\neq\bm{0}.$ (6)
@@
Let $\eta\colon\mathbb{R}^{3}\to\mathbb{R}$ be the radial extension of $f$ from Lemma 1, defined by $\eta(\bm{x}):=f(|\bm{x}|)$. Then $\hat{\eta}(\bm{\omega})=\hat{f}(|\bm{\omega}|)$, and thus $\hat{g}(\bm{\omega})=\hat{G}(\bm{\omega})\,\hat{\eta}(\bm{\omega})$ with $\hat{G}(\bm{\omega})=4\pi/|\bm{\omega}|^{2}$, i.e. $g=G*\eta$. See Appendix D.1 for details.
@@ 2.4 Window functions, periodization and Fourier coefficients
On the periodic domain $\Omega$, a periodized window function $\tilde{\varphi}$ is employed. More generally, for any function $f:\mathbb{R}^{3}\to\mathbb{R}$ we define the periodization $\tilde{f}$ on $\Omega$ by
@@
with coefficients (see Appendix D.2 for the derivation)
@@ 2.5 Prolate Spheroidal Wave Functions
The zeroth order [^1] Here “order” refers to the angular (azimuthal) index in the prolate spheroidal wave functions; order zero corresponds to the axisymmetric family. prolate spheroidal wave functions (PSWFs), $\{\psi_{j}^{c}\}_{j=0}^{\infty}$, arise as eigenfunctions of the integral operator
@@
$\lambda_{j}\psi_{j}^{c}(s)=\int_{-1}^{1}\psi_{j}^{c}(t)\,e^{\mathrm{i}cst}\,dt,\qquad s\in\mathbb{R},$ (23)
with eigenvalues ordered by decreasing magnitude
@@
The eigenfunctions $\psi_{j}^{c}$ are taken real and orthonormal in $L^{2}([-1,1])$, so that $\|\psi_{j}^{c}\|_{L^{2}([-1,1])}=1$. They alternate parity with $j$ (even for even $j$, odd for odd $j$) and are real–analytic on (1,1)(-1,1); in particular $\psi_{j}^{c}\in C^{\infty}([-1,1])$ [23]. For definiteness, we fix the sign of $\psi_{0}^{c}$ by requiring $\psi_{0}^{c}(0)>0$.
@@
Thus, in the present work we only need $\psi_{0}^{c}$, which is even has no zeros in (1,1)(-1,1) and, with the normalization $\psi_{0}^{c}(0)>0$, satisfies $\psi_{0}^{c}(s)>0$ for all $s\in(-1,1)$; in particular, $\psi_{0}^{c}(1)>0$ [23].
The Fourier transform of the truncation of $\psi_{0}^{c}$ to the interval [1,1][-1,1] satisfies a simple relation:
@@
Let $f:\mathbb{R}\to\mathbb{R}$ be defined by
$f(s)=\begin{cases}\psi_{0}^{c}(s),&s\in[-1,1],\\ 0,&\text{otherwise}.\end{cases}$ (27)
Then
$\hat{f}(\omega)=\lambda_{0}\,\psi_{0}^{c}(\omega/c),\qquad\omega\in\mathbb{R}.$ (28)
In particular,
$\hat{f}(\omega)=\lambda_{0}\,f(\omega/c),\qquad\omega\in[-c,c].$ (29)
Proof.
The Lemma follows directly from (23) letting $\omega=cs$. ∎
For computation, $\psi_{0}^{c}$ in [1,1][-1,1] can be obtained from its Legendre expansion and a tridiagonal eigenproblem derived from the Sturm–Liouville equation [24]
$(1-s^{2})\frac{\mathrm{d}^{2}\psi_{0}^{c}}{\mathrm{d}s^{2}}-2s\frac{\mathrm{d}\psi_{0}^{c}}{\mathrm{d}s}+(\chi_{0}-c^{2}s^{2})\psi_{0}^{c}=0,$ (30)
where $\chi_{0}>0$ is the differential-equation eigenvalue associated with $\psi^{c}_{0}$. Details of the computational problem are given in Appendix B.
The differential equation (30) is also useful for establishing several properties of the PSWFs. In order to do so, the following relation between the first eigenvalue $\chi_{0}$ and the bandlimit $c$ is useful.
Lemma 3.
Let $c>0$. Then the smallest eigenvalue associated with (30) satisfies
$\chi_{0}\leq\frac{c^{2}}{3}.$ (31)
From (31), it follows that for $s>1$, $c^{2}s^{2}-\chi_{0}\geq 2c^{2}/3>0$. This yields the following tail bound, which will be used in the error estimate in Proposition 1 of Section 5.1.
Lemma 4.
Let $c\geq\sqrt{3}$ and let $s\geq 1$. Then
$|\psi_{0}^{c}(s)|\leq\frac{\psi_{0}^{c}(1)}{s}.$ (32)
Proofs of Lemmas 3 and 4 are provided in Appendix C.
For $|s|>1$, $\psi_{0}^{c}$ exhibits oscillatory behavior, as described in the following lemma.
Lemma 5 (see [23, Theorem 3.44]).
Suppose that $c>0$ and that $s>1$. Suppose furthermore that $A^{c}(s)$ is defined via the formula
$A^{c}(s)=\frac{1}{\lambda_{0}\psi_{0}^{c}(1)}\int_{-1}^{1}\frac{\sin(c(s-t))\psi_{0}^{c}(t)t}{s-t}\,\mathrm{d}t.$ (33)
Then,
$\psi_{0}^{c}(s)=\frac{2\psi_{0}^{c}(1)}{cs\lambda_{0}}\left[\sin(cs)+A^{c}(s)\right].$ (34)
The representation (34) implies the approximation $\psi_{0}^{c}(s)\approx\frac{2\psi_{0}^{c}(1)}{\lambda_{0}}\frac{\sin(cs)}{cs}$, and the estimate $\psi_{0}^{c}(s)\lesssim\frac{2\psi_{0}^{c}(1)}{\lambda_{0}}\frac{1}{cs}$ for $s\geq 1$. These approximations are supported by numerical observations (see Figure 2) and are used in the error model in Section 5.1.1.
Figure 2: The tail of $|\psi_{0}^{c}(s)|$ for $c=\sqrt{3}$ and $c=10$, shown together with the bound (32), the approximation obtained from (34) by neglecting $A^{c}(s)$, and its corresponding upper bound. By Lemma 4, the bound $|\psi_{0}^{c}(s)|\leq\psi_{0}^{c}(1)/s$ holds for $c\geq\sqrt{3}$ (left). For large $s$, the approximation $\tfrac{2\psi_{0}^{c}(1)}{\lambda_{0}}\,\tfrac{|\sin(cs)|}{cs}$ accurately captures the oscillatory tail (right).
@@ 3 Ewald summation
Before we consider PSWF-based implementations, we briefly review the classical Ewald summation and its components. We then summarize the general fast Ewald algorithm, in which the PSWF will serve as both mollifier and window.
@@ 3.1 Ewald summation for triply periodic problems
Recall that the kernel split of a radial kernel $G\colon\mathbb{R}^{3}\to\mathbb{R}$ is
@@ 3.1.2 Direct computation of the Fourier-space sum
$\phi_{i}^{\text{far}}\approx\phi_{t}^{\text{far}}(\bm{x}_{i}):=\frac{1}{V}\sum_{\bm{k}\in\mathcal{I}_{\bm{m}}\setminus\{\bm{0}\}}\hat{M}(\bm{k})\,\hat{\rho}(\bm{k})\,e^{-\frac{2\pi}{L}\mathrm{i}\,\bm{k}\cdot\bm{x}_{i}},\qquad i=1,\ldots,n.$ (43)
Here, for the Laplace kernel $\hat{G}(\bm{\omega})=4\pi/|\bm{\omega}|^{2}$,
$\hat{M}(\bm{k})$ $=\hat{G}\bigl(\bm{\omega}(\bm{k})\bigr)\,\hat{\gamma}_{\sigma}\bigl(|\bm{\omega}(\bm{k})|\bigr)=\frac{L^{2}}{\pi|\bm{k}|^{2}}\,\hat{\gamma}_{\sigma}\bigl(|\bm{\omega}(\bm{k})|\bigr),$ (44)
@@
$\hat{\rho}(\bm{k})=\sum_{j=1}^{n}\rho_{j}\,e^{\tfrac{2\pi}{L}\mathrm{i}\,\bm{k}\cdot\bm{x}_{j}},\qquad\bm{k}\in\mathcal{I}_{\bm{m}}.$